# Pentagram maps over rings, Grassmannians, and skewers

### v2

@@ 6.2 Pseudo-difference operators
A (left) pseudo-difference operator over a ring $R$ is, basically, a formal Laurent series in terms of the left shift operator $T$ with coefficients in $R^{\mathbb{Z}}$. Formally, consider the subring $R^{\mathbb{Z}}[T,T^{-1}]$ of $\mathrm{End}_{R}(R^{\mathbb{Z}})$ generated by $R^{\mathbb{Z}}$, $T$, and $T^{-1}$. The ring $R^{\mathbb{Z}}[T,T^{-1}]$ consists of operators of the form $\sum\alpha_{i}T^{i}$, where $\alpha_{i}\in R^{\mathbb{Z}}$, and $\alpha_{i}\neq 0$ for only finitely many values of $i$. Consider a valuation-type function $\nu\colon R^{\mathbb{Z}}[T,T^{-1}]\to\mathbb{Z}_{+}\cup+\infty$ given by the degree of the lowest order term, where, by definition, $\nu(0)=+\infty$. Then $||\cdot||:=\exp(-\nu(\cdot))$ is a norm on $R^{\mathbb{Z}}[T,T^{-1}]$, in the sense of [9, Definition 2.11].
@@
The ring $R^{\mathbb{Z}}((T))$ of (left) pseudo-difference operators over a ring $R$ is the completion of $R^{\mathbb{Z}}[T,T^{-1}]$ with respect to the norm |||||$\cdot$||| | ⋅ | |.
@@
We have $(1+\mathcal{Q})^{-1}=\sum_{i=0}^{\infty}(-\mathcal{Q})^{i}$. The series converges since the ring $R^{\mathbb{Z}}((T))$ is complete, and $||\mathcal{Q}||\leq e^{-1}$. ∎
@@ 6.3 Lax representation for pentagram maps over rings
Given a projective equivalence class of polygons in $\mathbb{P}^{2}(R)$, consider the associated properly bounded difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{pb}$, and let $\mathcal{Q}:=\mathcal{D}_{+}^{-1}\mathcal{D}_{-}\in R^{\mathbb{Z}}((T))$ (this is well-defined by Corollary 6.8). Note that since $\mathcal{D}$ is defined up to the left-right action of $(R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}$, the operator $\mathcal{Q}$ is defined up to the conjugation action of $(R^{\mathbb{Z}})^{*}$ on $R^{\mathbb{Z}}((T))$. Hence, we have a map $\mathfrak{P}(\mathbb{P}^{2}(R))\to R^{\mathbb{Z}}((T))/(R^{\mathbb{Z}})^{*},$ taking a projective equivalence class of polygons to the conjugacy class of the associated operator $\mathcal{D}_{+}^{-1}\mathcal{D}_{-}$. We claim that this map provides a Lax representation of the pentagram map in the sense that polygons related by the pentagram map are being mapped to conjugate elements of $R^{\mathbb{Z}}((T))/(R^{\mathbb{Z}})^{*}$.
@@
Suppose $(P_{i})$ is a polygon in $\mathbb{P}^{2}(R)$ whose image $(\tilde{P}_{i})$ is also a polygon. Then the associated operators $\mathcal{Q}=\mathcal{D}_{+}^{-1}\mathcal{D}_{-}$ and $\tilde{\mathcal{Q}}=\tilde{\mathcal{D}}_{+}^{-1}\tilde{\mathcal{D}}_{-}$ (defined up to the conjugation action of $(R^{\mathbb{Z}})^{*}$) are conjugate in $R^{\mathbb{Z}}((T))$.
@@
By Theorem 6.5, the operators $\mathcal{D}$ and $\tilde{\mathcal{D}}$ can be chosen so that
$\mathcal{D}_{+}\tilde{\mathcal{D}}_{-}=\mathcal{D}_{-}\tilde{\mathcal{D}}_{+}.$
But then
$\tilde{\mathcal{Q}}=\tilde{\mathcal{D}}_{+}^{-1}\tilde{\mathcal{D}}_{-}=\tilde{\mathcal{D}}_{+}^{-1}\mathcal{Q}\tilde{\mathcal{D}}_{+},$ (6.5)
as desired. ∎
We note that, when $R$ is a field, (6.5) coincides with the Lax representation of the classical pentagram map given in [6, 8].
@@ 7 Invariants of pentagram maps on closed polygons
Here we use the Lax form found in the previous section to give invariants (i.e., conserved quantities, or first integrals) of the pentagram map on closed polygons (i.e., polygons given by periodic sequences $P_{i}\in\mathbb{P}^{2}(R)$). In the case of a field, these invariants are known to form a maximal Poisson-commuting family [15]. We expect that similar results can be obtained for much bigger classes of rings.
@@ 7.1 Closed polygons and super-periodic difference operators
Let $\mathcal{D}$ be a super-periodic left properly bounded difference operator with period $n$. Consider also $\alpha,\beta\in(\mathbb{R}^{\mathbb{Z}})^{*}$. Then, $\alpha\mathcal{D}\beta$ is super-periodic with period $n$ if and only if $\alpha$ and $\beta$ are $n$-periodic.
@@
Conversely, suppose that that $\alpha\mathcal{D}\beta$ is super-periodic with period $n$. Fix $i\in\mathbb{Z}$, and choose $a\in\mathrm{Ker}\,\alpha\mathcal{D}\beta$ such that $a_{i}=1$. Since $\alpha\mathcal{D}\beta$ is super-periodic with period $n$, we also have $a_{i+n}=1$. Further, since $a\in\mathrm{Ker}\,\alpha\mathcal{D}\beta$, we have $\beta a\in\mathrm{Ker}\,\,\mathcal{D}$. Since $\mathcal{D}$ is super-periodic with period $n$, this gives $\beta_{i}a_{i}=\beta_{i+n}a_{i+n}$, i.e, $\beta_{i+n}=\beta_{i}$. Since $i$ was arbitrary, this means that $\beta$ is $n$-periodic. But then $\alpha$ is a product of $n$-periodic pseudo-difference operators:
@@
If $\mathcal{D}$ is a super-periodic operator with period $n$, then $\varepsilon_{\mathcal{D}}$ is an $n$-periodic sequence in $(\mathrm{Ker}\,\mathcal{D})^{*}$. So, the mapping $\bar{\phi}\colon R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ takes super-periodic operators with period $n$ to closed polygons with the same period. Conversely, given a closed $n$-gon $(P_{i})$, we can choose spanning vectors $v_{i}\in P_{i}$ so that $v_{i+n}=v_{i}$, and then the operator $\mathcal{D}$ such that $\mathcal{D}v=0$ can be taken to be $n$-periodic. So, the set of properly bounded super-periodic left difference operators of degree $d$ and period $n$ is mapped onto the set of projective equivalence classes of closed $n$-gons . Furthermore, by Proposition 7.2 preimages of every class of polygons under this mapping are precisely the orbits of the left-right action of $n$-periodic sequences of units.∎
@@ 7.2 The trace of a periodic pseudo-difference operator
For a ring $R$, consider the additive subgroup $[R,R]\subset R$ generated over $\mathbb{Z}$ by all commutators $[a,b]:=ab-ba$. The Abelian group $R/[R,R]$ is known as the cyclic space of $R$, and the quotient map $R\to R/[R,R]$ as the universal trace.
@@
Thus, the trace is a homomorphism of Abelian groups $\mathrm{Tr}\colon R^{\mathbb{Z}}((T))\to R/[R,R]$. The following shows it is indeed a trace (i.e., satisfies the cyclic property).
@@
We need to show that $\mathrm{Tr}(\mathcal{Q}\widehat{\mathcal{Q}}-\widehat{\mathcal{Q}}\mathcal{Q})=0$. Note that the left-hand side is bilinear over $\mathbb{Z}$. Furthermore, if we endow the cyclic space with discrete topology, then the trace function becomes continuous, so the operation $\mathcal{Q},\widehat{\mathcal{Q}}\mapsto\mathrm{Tr}(\mathcal{Q}\widehat{\mathcal{Q}}-\widehat{\mathcal{Q}}\mathcal{Q})$ distributes over infinite sums. Hence, it suffices to verify the statement in the case $\mathcal{Q}=\alpha T^{k}$, $\widehat{\mathcal{Q}}=\beta T^{{l}}$, where $\alpha,\beta\in R^{\mathbb{Z}}$ are $n$-periodic. Clearly, $\mathrm{Tr}\,\mathcal{Q}\widehat{\mathcal{Q}}=\mathrm{Tr}\,\widehat{\mathcal{Q}}\mathcal{Q}=0$ if $l\neq-k$, so suppose $l=-k$. Then
@@ 7.3 Invariants of the pentagram map
As discussed above, to any projective equivalence class of closed $n$-gons we can assign an $n$-periodic pseudo-difference operator $\mathcal{Q}:=\mathcal{D}_{+}^{-1}\mathcal{D}_{-}$, defined up to conjugation by $n$-periodic bi-infinite sequences of units. Upon application of the pentagram map, the operator $\mathcal{Q}$ changes to a conjugate one. Now, consider the functions on the space $\mathfrak{P}(\mathbb{P}^{2}(R))$ of projective equivalence classes of polygons defined by
$f_{ij}:=\mathrm{Tr}(T^{in}\mathcal{\mathcal{Q}}^{j}),\hskip 28.45274pti\in\mathbb{Z},j\in\mathbb{Z}_{\geq 0},$ (7.1)
where $\mathcal{Q}$ is defined as described above.
@@
    The functions $f_{ij}$ are independent of the choice of $\mathcal{Q}$ and, hence, are well-defined functions on $\mathfrak{P}(\mathbb{P}^{2}(R))$.
@@
Since $\mathcal{Q}$ is defined up to conjugation, and the action of the pentagram map is also given by conjugation, it suffices to show that the functions $\mathcal{Q}\mapsto\mathrm{Tr}(T^{in}\mathcal{\mathcal{Q}}^{j})$ are conjugation-invariant (central). Indeed,
@@
The ring of $n$-periodic pseudo-difference operators over $R$ is isomorphic to the ring $\text{M}_{n}(R)((z))$ of formal Laurent series over $\text{M}_{n}(R)$, cf. [8, Remark 3.8]. The isomorphism is given by
@@
Here $E_{i,j}\in\text{M}_{n}(R)$ is the matrix with a $1$ at position $(i,j)$ and zeros elsewhere. Therefore, in the periodic case, Lax representation (6.5) can be thought of as being valued in $\text{M}_{n}(R)((z))$. In other words, it is a Lax representation with spectral parameter. Also note that, for $R=\text{M}_{k}(\mathbb{F})$, we have $\text{M}_{n}(R)\simeq\text{M}_{nk}(\mathbb{F})$. So, in this case, one gets Lax matrices whose entries are in a field, allowing one to study the corresponding (Grassmann) pentagram maps using standard tools of integrable systems theory. This Lax representation is different from and in a sense dual to the one given in [4].

### v3

@@ 6.2 Pseudo-difference operators
A (left) pseudo-difference operator over a ring $R$ is a formal Laurent series in terms of the left shift operator $T$ with coefficients in $R^{\mathbb{Z}}$. Formally, consider the subring $R^{\mathbb{Z}}[T,T^{-1}]$ of $\mathrm{End}_{R}(R^{\mathbb{Z}})$ generated by $R^{\mathbb{Z}}$, $T$, and $T^{-1}$. The ring $R^{\mathbb{Z}}[T,T^{-1}]$ consists of operators of the form $\sum\alpha_{i}T^{i}$, where $\alpha_{i}\in R^{\mathbb{Z}}$, and $\alpha_{i}\neq 0$ for only finitely many values of $i$. Consider a valuation-type function $\nu\colon R^{\mathbb{Z}}[T,T^{-1}]\to\mathbb{Z}_{+}\cup+\infty$ given by the degree of the lowest order term, where, by definition, $\nu(0)=+\infty$. Then $||\cdot||:=\exp(-\nu(\cdot))$ is a norm on $R^{\mathbb{Z}}[T,T^{-1}]$, in the sense of [10, Definition 2.11].
@@
The ring $R^{\mathbb{Z}}((T))$ of (left) pseudo-difference operators over a ring $R$ is the completion of $R^{\mathbb{Z}}[T,T^{-1}]$ with respect to the norm $||\cdot||$.
@@
We have $(1+\mathcal{Q})^{-1}=\sum_{i=0}^{\infty}(-\mathcal{Q})^{i}$. The series converges since the ring $R^{\mathbb{Z}}((T))$ is complete, and $\mathcal{Q}||\leq e^{-1}$. ∎
@@ 6.3 Lax representation for pentagram maps over rings
In this section, we recall the notion of a Lax pair and construct such pairs for pentagram maps over rings. Details on Lax pairs can be found in any textbook on integrable systems, see, e.g., [3, Section 2.4].
Let $X$ be a set, and $\phi\colon X\dashrightarrow X$ be a dynamical system. We use a dashed arrow to emphasize that the map $\phi$ might not be everywhere defined. For instance, the pentagram map is not defined on the whole set of polygons, since the image of a polygon under that map might not be a polygon.
Let $S$ be a ring, and $S^{\times}$ be its group of units. A Lax pair for $\phi$ is a pair of maps $L\colon X\dashrightarrow S$, $A\colon X\dashrightarrow S^{\times}$, defined on the domain of $\phi$, such that
$L(\phi(x))=A(x)^{-1}L(x)A(x)$
whenever $\phi(x)$ is in the domain of $\phi$. To distinguish meaningful Lax pairs from trivial cases such as $L\equiv 0$, $A\equiv 1$, we say that a Lax pair is faithful if the mapping $L$ is injective.
An immediate consequence of the existence of a Lax pair is that functions of the form $f(L(x))$, where $f$ is central (i.e., invariant under conjugation by units) on $S$, are invariants of $\phi$. In many examples, it is possible to show that, for a suitable Poisson bracket on $X$, such invariants constitute a maximal Poisson-commutative family, implying that $\phi$ is an Arnold-Liouville integrable system.
Another framework where Lax pairs play a central role is algebraic integrability. In that setting, $S$ is supposed to be the ring of matrices depending on an additional parameter, known as the spectral parameter. In this case, the characteristic polynomial of $\det(L(x)-\lambda\mathbf{1})$ is, for fixed $x$, a function of two variables, $\lambda$ and the spectral parameter. The zero locus of that function is a plane curve, known as the spectral curve. The spectral curve is preserved by the dynamical system $\phi$. Moreover, in many cases, one can show that, for a fixed spectral curve, the dynamics of $\phi$ can be identified with linear motion on the Jacobian of the spectral curve, turning $\phi$ into an algebraic completely integrable system.
The pentagram map in $\mathbb{P}^{2}(\mathbb{F})$, where $\mathbb{F}$ is $\mathbb{R}$ or $\mathbb{C}$, is Arnold-Liouville integrable [18]; it is also algebraically integrable for any field $\mathbb{F}$ with $\mathrm{char}\,\mathbb{F}\neq 2$ [22, 30]. A Lax pair with spectral parameter is given, e.g., in [7, Proposition 4.10].
In the setting of rings, one cannot expect neither Arnold-Liouville, nor algebraic integrability without further assumptions. Indeed, neither Poisson brackets, nor algebraic curves and their Jacobians make sense in this generality. So, we take the existence of a Lax pair as a definition of integrability, cf. [5, 12] . In the case of closed polygons, our Lax pair possesses a spectral parameter, see Remark 7.9. We expect that, under further assumptions, our Lax pair will yield both Arnold-Liouville and algebraic integrability, since over the real and complex numbers it specializes to the standard Lax pair which is known to possess these properties.
The construction of a Lax pair for the pentagram map over a stably finite ring $R$ is as follows. Denote the pentagram map by $\Phi$ and let $\mathcal{P}\in\mathfrak{P}(\mathbb{P}^{2}(R))$ be a projective equivalence class of polygons in $\mathbb{P}^{2}(R)$ which is in the domain of $\Phi$ (i.e., the images of polygons in $\mathcal{P}$ under the pentagram map are again polygons). Then $\Phi(\mathcal{P})\in\mathfrak{P}(\mathbb{P}^{2}(R))^{\mathrm{reg}}$. Let $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{reg}}$ be an operator associated to $\Phi\left(\mathcal{P}\right)$. Then, by Theorem 6.5, the equation
$\mathcal{D}_{+}\widehat{\mathcal{D}}_{-}=\mathcal{D}_{-}\widehat{\mathcal{D}}_{+}$ (6.7)
has a solution $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$, unique up to the left action of $(R^{\mathbb{Z}})^{\times}$, such that the associated class of polygons coincides with $\mathcal{P}$ (note that we have switched the roles of $\mathcal{D}$ and $\widehat{\mathcal{D}}$). Furthermore, we know that $\mathcal{D}_{+}$ is a unit in $R^{\mathbb{Z}}((T))$ by Corollary 6.8. Set
$\mathcal{L}(\mathcal{P}):=\mathcal{D}_{+}^{{-1}}\mathcal{D}_{-},\quad\mathcal{A}(\mathcal{P}):=\widehat{\mathcal{D}}_{+}$. (6.8)
Note that the definition of $\mathcal{L}(\mathcal{P})$ is independent on the choice of solution of equation (6.7). However, both $\mathcal{L}(\mathcal{P})$ and $\mathcal{A}(\mathcal{P})$ depend on the choice of $\widehat{\mathcal{D}}$. Upon replacing $\widehat{\mathcal{D}}$ by $\alpha\widehat{\mathcal{D}}\beta$, where $\alpha,\beta\in(R^{\mathbb{Z}})^{\times}$, these operators transform as
$\mathcal{L}(\mathcal{P})\mapsto\alpha\mathcal{L}(\mathcal{P})\alpha^{-1},\quad\mathcal{A}(\mathcal{P})\mapsto\alpha\mathcal{A}(\mathcal{P})\beta.$ (6.9)
@@
The maps $\mathcal{L},\mathcal{A}$, defined by (6.8) up to the action (6.9), constitute a faithful Lax pair for the pentagram map in $\mathbb{P}^{2}(R)$.
@@
From (6.7), we get
$\mathcal{L}(\mathcal{P})=\mathcal{D}_{+}^{-1}\mathcal{D}_{-}=\widehat{\mathcal{D}}_{-}\widehat{\mathcal{D}}_{+}^{-1}.$ (6.10)
So,
$\mathcal{A}(\mathcal{P})^{-1}\mathcal{L}(\mathcal{P})\mathcal{A}(\mathcal{P})=\widehat{\mathcal{D}}_{+}^{-1}\widehat{\mathcal{D}}_{-}$,
which, up to conjugation action of $(R^{\mathbb{Z}})^{\times}$, is equal to $\mathcal{L}(\Phi(\mathcal{P}))$, as needed. To prove that this Lax pair is faithful, we need to show that if
$\mathcal{D}_{+}^{-1}\mathcal{D}_{-}=\tilde{\mathcal{D}}_{+}^{-1}\tilde{\mathcal{D}}_{-}$ (6.11)
for some $\tilde{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$, then $\tilde{\mathcal{D}}$ represents the same class of polygons $\mathcal{P}$ as $\mathcal{D}$. Indeed, from (6.10) and (6.11), we get
$\tilde{\mathcal{D}}_{+}^{-1}\tilde{\mathcal{D}}_{-}=\widehat{\mathcal{D}}_{-}\widehat{\mathcal{D}}_{+}^{-1},$
so
$\tilde{\mathcal{D}}_{+}\widehat{\mathcal{D}}_{-}=\tilde{\mathcal{D}}_{-}\widehat{\mathcal{D}}_{+}$.
By Theorem 6.5, the latter means that the class of polygons represented by $\tilde{\mathcal{D}}$ is the image of $\Phi(\mathcal{P})$ under the inverse pentagram map, i.e., $\mathcal{P}$.∎
We note that, when $R$ is a field, this Lax representation becomes the one given in [7, 9].
@@ 7 Invariants of pentagram maps on closed polygons
Here we use the Lax form found in the previous section to give invariants (i.e., conserved quantities, or first integrals) of the pentagram map on closed polygons (i.e., polygons given by periodic sequences $P_{i}\in\mathbb{P}^{2}(R)$). In the case of the pentagram map over real numbers, these invariants (in a slightly more general setting of twisted polygons) are known to form a maximal Poisson-commuting family [18, Theorem 1]. We expect that similar results can be obtained for much bigger classes of rings.
Throughout this section, the ground ring $R$ is assumed to be stably finite.
@@ 7.1 Closed polygons and super-periodic difference operators
Let $\mathcal{D}$ be a super-periodic left properly bounded difference operator with period $n$. Consider also $\alpha,\beta\in(R^{\mathbb{Z}})^{\times}$. Then, $\alpha\mathcal{D}\beta$ is super-periodic with period $n$ if and only if $\alpha$ and $\beta$ are $n$-periodic.
@@
Conversely, suppose that that $\alpha\mathcal{D}\beta$ is super-periodic with period $n$. Fix $i\in\mathbb{Z}$, and choose $a\in\mathrm{Ker}\,\alpha\mathcal{D}\beta$ such that $a_{i}=1$. Since $\alpha\mathcal{D}\beta$ is super-periodic with period $n$, we also have $a_{i+n}=1$. Further, since $a\in\mathrm{Ker}\,\alpha\mathcal{D}\beta$, we have $\beta a\in\mathrm{Ker}\,\,\mathcal{D}$. Since $\mathcal{D}$ is super-periodic with period $n$, this gives $\beta_{i}a_{i}=\beta_{i+n}a_{i+n}$, so, $\beta_{i+n}=\beta_{i}$. Additionally, given that $i$ was arbitrary, this means that $\beta$ is $n$-periodic. But then $\alpha$ is a product of $n$-periodic pseudo-difference operators:
@@
If $\mathcal{D}$ is a super-periodic operator with period $n$, then $\varepsilon_{\mathcal{D}}$ is an $n$-periodic sequence in $(\mathrm{Ker}\,\mathcal{D})^{*}$. So, the mapping $\bar{\phi}\colon R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ takes super-periodic operators with period $n$ to closed polygons with the same period. Conversely, given a closed $n$-gon $(P_{i})$, we can choose spanning vectors $v_{i}\in P_{i}$ so that $v_{i+n}=v_{i}$, and then the operator $\mathcal{D}$ such that $\mathcal{D}v=0$ can be taken to be $n$-periodic. So, the set of properly bounded super-periodic left difference operators of degree $d$ and period $n$ is mapped onto the set of projective equivalence classes of closed $n$-gons . Furthermore, by Proposition 7.2 preimages of every class of polygons under this mapping are precisely the orbits of the left-right action of $n$-periodic sequences of units.∎
@@ 7.2 The trace of a periodic pseudo-difference operator
For a ring $R$, consider the additive subgroup $R,R]\subset R$ generated over $\mathbb{Z}$ by all commutators [a,b]:=abba[a,b]:=ab-ba. The Abelian group $R/[R,R]$ is known as the cyclic space of $R$, and the quotient map $R\to R/[R,R]$ as the universal trace.
@@
Thus, the trace is a homomorphism of Abelian groups $\mathrm{Tr}\colon R^{\mathbb{Z}}((T))\to R/[R,R]$. The following shows it is indeed a trace (that is, it satisfies the cyclic property).
@@
We need to show that $\mathrm{Tr}(\mathcal{Q}\widehat{\mathcal{Q}}-\widehat{\mathcal{Q}}\mathcal{Q})=0$. Note that the left-hand side is bilinear over $\mathbb{Z}$. Furthermore, if we endow the cyclic space with discrete topology, then the trace function becomes continuous, so the operation $\mathcal{Q},\widehat{\mathcal{Q}}\mapsto\mathrm{Tr}(\mathcal{Q}\widehat{\mathcal{Q}}-\widehat{\mathcal{Q}}\mathcal{Q})$ distributes over infinite sums. Hence, it suffices to verify the statement in the case $\mathcal{Q}=\alpha T^{k}$, $\widehat{\mathcal{Q}}=\beta T^{{l}}$, where $\alpha,\beta\in R^{\mathbb{Z}}$ are $n$-periodic. Clearly, $\mathrm{Tr}\,\mathcal{Q}\widehat{\mathcal{Q}}=\mathrm{Tr}\,\widehat{\mathcal{Q}}\mathcal{Q}=0$ if $l\neq-k$, so suppose l=kl=-k. Then
@@ 7.3 Invariants of the pentagram map
By the above, to any projective equivalence class $\mathcal{P}$ of closed $n$-gons we can assign an $n$-periodic pseudo-difference operator $\mathcal{L}(\mathcal{P}):=\mathcal{D}_{+}^{-1}\mathcal{D}_{-}$, defined up to conjugation by $n$-periodic bi-infinite sequences of units. Upon application of the pentagram map, the operator $\mathcal{L}(\mathcal{P})$ changes to a conjugate one. Now, consider the functions on the space $\mathfrak{P}(\mathbb{P}^{2}(R))$ of projective equivalence classes of polygons defined by
$f_{ij}(\mathcal{P}):=\mathrm{Tr}(T^{in}\mathcal{L}(\mathcal{P})^{j}),\hskip 28.45274pti\in\mathbb{Z},j\in\mathbb{Z}_{\geq 0},$ (7.1)
where $\mathcal{L}$ is defined as described above.
@@
    The functions $f_{ij}$ are independent of the choice of $\mathcal{L}(\mathcal{P})$ and, hence, are well-defined functions on $\mathfrak{P}(\mathbb{P}^{2}(R))$ valued in the cyclic space $R/[R,R]$.
@@
Since $\mathcal{L}(\mathcal{P})$ is defined up to conjugation, and the action of the pentagram map is also given by conjugation, it suffices to show that the functions $\mathcal{Q}\mapsto\mathrm{Tr}(T^{in}\mathcal{\mathcal{Q}}^{j})$ on $n$-periodic pseudo-difference operators are conjugation-invariant (central). Indeed,
@@
Example 7.7
If $R=\mathrm{M}_{k}(\mathbb{F})$, then $R/[R,R]\simeq\mathbb{F}$, so the invariants $f_{ij}$ are valued in the ground field $\mathbb{F}$.
Example 7.8
The ring $R=\mathbb{D}$ of dual numbers is commutative, so $[R,R]=0$, and the invariants $f_{ij}$ are valued in the ring $\mathbb{D}$ itself. The real parts of these invariants can be understood as follows. There is a natural projection $\pi\colon\mathbb{P}^{2}(\mathbb{D})\to\mathbb{P}^{2}(\mathbb{R})$ which sends every affine line to its direction. The skewer pentagram map and the ordinary pentagram map are semi-conjugate via this projection (or, more precisely, via the induced projection on polygon spaces). In particular, $\pi$-lifts of the invariants of the usual pentagram map are invariants of the skewer map. These $\pi$-lifts are precisely the real parts of $f_{ij}$ (this easily follows from the fact that the real part function $\Re\colon\mathbb{D}\to\mathbb{R}$ is a homomorphism of rings and hence takes all objects associated with the skewer map to the corresponding objects for the pentagram map over the reals). The non-real parts of $f_{ij}$ are additional invariants which do not come from the pentagram map in $\mathbb{P}^{2}(\mathbb{R})$.
@@
The ring of $n$-periodic pseudo-difference operators over $R$ is isomorphic to the ring $\mathrm{M}_{n}(R)((z))$ of formal Laurent series over $\mathrm{M}_{n}(R)$, cf. [9, Remark 3.8]. The isomorphism is given by
@@
Here $E_{i,j}\in\mathrm{M}_{n}(R)$ is the matrix with a $1$ at position $(i,j)$ and zeros elsewhere. Therefore, in the periodic case, the Lax representation given by Proposition 6.10 can be thought of as being valued in $\mathrm{M}_{n}(R)((z))$. In other words, it is a Lax representation with spectral parameter. Also note that, for $R=\mathrm{M}_{k}(\mathbb{F})$, we have $\mathrm{M}_{n}(R)\simeq\mathrm{M}_{nk}(\mathbb{F})$. So, in this case, one gets Lax matrices whose entries are in a field, allowing one to study the corresponding (Grassmann) pentagram maps using standard tools of integrable systems theory. This Lax representation is different from and in a sense dual to the one given in [5].